# Modeling Advection-Dominated Flows with Space-Local Reduced-Order Models

### v2

@@ 5.2 Projection error
For the construction of the L-POD and LO-POD ROMs we have to determine the number of subdomains $I$ and the number of modes per subdomain $q$, with $Iq=r$, which results in a basis that generalizes best outside the training data. To do this we evaluate the projection error, see (41), on both the training and validation data. The results are depicted in Figure 5 for different $I$, averaged over both the training (solid line) and validation data (dashed line).
Figure 5: (Top) Projection error evaluated over the training (solid line) and validation (dashed line) data for different $I$ for the space-local approaches. (Bottom) Projection error for $I=10$ for L-POD and LO-POD evaluated over the training (solid line) and validation (dashed line) data. The projection error for G-POD is also depicted.
A basis that generalizes well should perform well on both the training and validation data. We observe that for $I=5$ the training error is lowest, but the validation error is rather large, so the basis does not generalize well. For higher values of $I$, the training and validation errors are much closer, but tend to increase with increasing $I$. In this case, $I=10$ is considered optimal as the training and validation error are small and similar in size. This is true for both L-POD and LO-POD. For the remainder of this text we will therefore stick to $I=10$ for the construction of the local ROMs. In general, the choice for $I$ is likely problem dependent and an a-priori study similar to Figure 5 needs to be performed to choose an optimal value.
Next, we compare the performance of the space-local approaches against each other, as well as to G-POD. This is also displayed in Figure 5 (bottom plot). We observe that the projection error converges faster for LO-POD, as compared to L-POD. This can be explained by the fact the LO-POD basis functions are constructed from twice as many finite difference points as for L-POD, due to the overlapping subdomains. This means the fit is carried out over twice as many DOF, which yields a lower projection error. The difference between convergence on the training and validation data is small for both space-local approaches. On the other hand, for G-POD the convergence of the projection error on the training data is very fast, see (13), but on the validation data the error hardly converges. This is because the G-POD is only suited for representing the wave in the left side of the domain, as detailed in the next section. We conclude that the space-local approaches result in a basis that generalizes better than the global approach, for this particular problem.
@@ 5.3 Resulting basis
Figure 6: The first six G-POD modes and the first six space-local POD modes for L-POD and LO-POD. For visualization purposes the space-local POD modes are displayed on adjacent subdomains.
We find that the G-POD modes are only nonzero where the traveling wave is represented in the training data. This explains the fact that the error on the validation data hardly converged in Figure 5. For the space-local approaches, where a common basis is “copy/pasted” across the entire domain, this is not an issue. Regarding L-POD we observe that the obtained basis functions do not smoothly decay to zero at the edge of the subdomain, but instead end abruptly. Finally, we observe that for LO-POD the post-processing procedure in (28) indeed results in a basis that smoothly decays on the edge of the subdomains.
@@ 5.4 Accuracy and energy conservation of ROM
Having obtained a basis for each of the POD approaches we construct a set of ROMs. Based on the results in section 5.2 we select $r=60$, $I=10$, and $q=6$ for the space-local approaches. To keep the size of the basis the same we select $r=60$ for G-POD. The ROM results are obtained by evaluating (35) from the initial condition in (40). The resulting simulations up to $t=5$ are presented in Figure 7.
@@
For G-POD we find that after the training data the performance degrades significantly. This exposes the limitations of G-POD, as it is not able to adapt to the wave traveling outside the training range. The space-local approaches perform better in this regard. Both L-POD and LO-POD are capable of extrapolating the traveling of the wave past the training region. However, L-POD seems to suffer from discontinuities in $\mathbf{u}_{r}^{\text{ROM}}$, as the edges of the subdomains become increasingly visible as the simulation progresses. LO-POD does not suffer from this issue and smoothly extrapolates the solution past the training region. This means the smooth reconstruction coming from LO-POD indeed improves the quality of the simulation for the same number of DOF.
@@
The results are depicted for both the FOM time step size $\Delta t$ (solid line) and a five times larger time step $5\cdot\Delta t$ (dashed line). This larger time step is based on an eigenvalue analysis of the ROM operator, also presented in Figure 8. For this analysis we determined the eigenvalues of the operator $\mathbf{D}$ projected on the ROM basis. This operator is given by $\mathbf{P}\mathbf{D}\mathbf{P}$. It can be shown that replacing $\mathbf{D}$ in the FOM, see (6), by this projected operator is equivalent to integrating the ROM, see (35). The eigenvalues $\lambda_{i}$ are ordered according to the magnitude of their absolute values. We observe that the largest eigenvalue for the space-local approaches is roughly five times smaller than the one of G-POD and the FOM. This is likely caused by the fact that the G-POD basis functions contain higher frequencies than the space-local basis functions, see Figure 6. Using a smaller G-POD basis could alleviate this. Based on this analysis we also evaluate the performance of the ROMs for a five times larger time step [35].
@@
Looking at the solution error we observe that G-POD performs best in the training region (the error is so small that it is outside the plotting range). However, after leaving the training region ($t>1$) the performance degrades rapidly. On the other hand, for the space-local approaches we do not observe this jump of error outside the training region, but rather a steady increase. Importantly, outside the training region the space-local approaches are much more accurate than G-POD. In particular LO-POD improves with more than one order of magnitude upon both G-POD and L-POD outside the training region.
@@
When increasing the size of the time step by five-fold the simulation quickly becomes unstable for G-POD. For L-POD both time steps yield stable simulations, giving results that are very close to each other. For LO-POD and a smaller time step size the error seems to converge to an equilibrium. However, for a larger time step it increases steadily. This means there is likely still a benefit to taking a smaller time step. An interesting continuation of this research would be to find a way to systematically determine the “sweet spot” between increasing the time step and maintaining accuracy of the ROM.
@@
In terms of energy conservation, we observe that all ROMs conserve the energy as predicted by the theoretical analysis, except for a time discretization error incurred by the use of RK4. This error increases when the time step size is increased. Only for G-POD with an increased time step size the simulation becomes unstable. For LO-POD the numerical error in the instantaneous change in energy, see (39), is the largest. This is likely due the linear system that needs to be solved to evaluate the ROM for a non-orthogonal basis, see (35). However, the time discretization error is the main source of error, as the change in energy during the simulation is roughly the same for L-POD and LO-POD.
@@ 5.5 Convergence with increasing ROM dimension
Next, we look at the convergence of the solution error as we increase the size of ROM basis $r$. In addition, we compare the ROMs to a finite difference discretization with the same number of DOF. Note that for the ROMs $\text{DOF}=r$. For the finite difference simulations we first project the solution on the FOM grid using linear interpolation. We then compute the difference, such as in (41), and compute the $L_{2}$-norm to quantify the error. For the local ROMs we stick to $I=10$ subdomains, while increasing the number of basis functions per subdomain $q$ to increase $r$. We consider a maximum of $r=100$ for the ROMs.
@@
Regarding the local ROMs we observe rapid convergence of the solution error as we increase the number of DOF, with LO-POD consistently outperforming L-POD. This is in line with the projection error convergence discussed in section 5.2. Regarding G-POD we observe no convergence of the solution error. Regarding the finite difference discretization we find it converges at a much slower rate than the space-local ROMs.
To obtain a conclusive answer about the scaling of the ROMs we considered the number of nonzero entries in the ROM operator $\mathbf{A}$. This is also depicted in Figure 9. Here we find that the number of nonzero entries scales according to a power law, with L-POD scaling the most favorably. This is in line with the discussion presented in section 4.2. Hyperreduction can be carried out to further reduce the computational cost of the ROMs [32, 33, 17]. Finally, one could also increase the time step size, as discussed in section 5.4, to further decrease the computational cost of the space-local ROMs.
5.6 2D Advection equation
To further evaluate the viability of the space-local ROMs we consider the 2D advection equation:
@@
where some quantity $u(\mathbf{x},t)\in\mathbb{R}$ is advected through space $\mathbf{x}\in\mathbb{R}^{2}$ by a stationary velocity field $\mathbf{V}\in\mathbb{R}^{2}$ which is divergence free, i.e. $\nabla\cdot\mathbf{V}=0$. The fact that the velocity field is now space-dependent, as opposed to the 1D case where it was uniform, adds to the difficulty of the test case. The diffusion rate is determined by the scalar viscosity $\nu\geq 0$. In our case we employ a finite volume discretization with a staggered grid for the velocity field [36]. This is done to preserve the skew-symmetry of the operator. The diffusive term is discretized using a simple second order scheme.
For the FOM we consider the following setting: A periodic domain $\Omega=[-\pi,\pi]\times[-\pi,\pi]$, a velocity field given by $\mathbf{V}=[\cos(x-y),\cos(x-y)]^{T}$ discretized on a $256\times 256$ uniform grid, and a viscosity of $\nu=10^{-3}$. The initial condition is given by the sum of three Gaussians, each centered on one of the streamlines of the flow where the velocity is highest:
$\begin{split}u(\mathbf{x},0)=&-\exp(-2x^{2}-2y^{2})+\exp(-2(x-\frac{\pi}{2})^{2}-2(y+\frac{\pi}{2})^{2})\\ &+\exp(-2(x+\frac{\pi}{2})^{2}-2(y-\frac{\pi}{2})^{2}).\end{split}$ (43)
The FOM is integrated in time using a RK4 scheme with $\Delta t=0.025$. Each 16 time steps we save a snapshot of $u(x,t)$ for the construction of the ROM basis. The snapshots collected in the interval $t=[0,12]$ are used as training data, the interval $t=(16,20]$ is used as validation data, and the interval $t=(20,40]$ is used to test the extrapolation capabilities of the ROMs. The ROMs are also integrated in time explicitly using a RK4 scheme. However, for the LO-POD ROM we resort to the implicit Crank-Nicolson method, see [24], to limit the computational cost. This scheme allows us to obtain second order accuracy in time, while only solving a single linear system each time step, instead of each evaluation of (35) in the RK4 scheme. To solve the resulting system efficiently we employ an LU-decomposition [37]. This is required, as the LO-POD basis is non-orthogonal, i.e. $\mathbf{S}\neq\mathbf{I}$.
Regarding the subdomain decomposition we use the fact that the FOM is discretized on a uniform grid. In this way we can simply subdivide the domain uniformly into subregions. For L-POD this results in a straightforward decomposition. For LO-POD the decomposition is chosen such that each subdomain overlaps with its 4 neighboring subdomains which share a vertex in the middle of the considered domain. This is depicted in Figure 10.
@@
The number of subdomains $I$ for the space-local approaches and the number of modes $q$ is chosen according to the performance on the validation set. This is presented in B. Based on this we settle on $I=8\times 8$ and $q=20$ for L-POD and $q=15$ for LO-POD. We settle on these values for $q$ as these are the lowest values which surpass a projection error threshold of $10^{-2}$ on the validation set for this value of $I$. The value of $I$ is chosen as it generalizes well from the training data set to the validation data set. For G-POD we choose $r=q=31$, as the snapshot matrix contains $31$ snapshots for G-POD such that the SVD only produces $31$ basis functions. For the size of the time step we choose the largest value which does not degrade the ROM performance. This is presented in C. The time step sizes used are presented in Table 1.
@@
The resulting basis computed with the SVD for each of the POD approaches are depicted in Figure 11.
@@
From the basis functions we can see that the flow runs diagonally across the domain. This is expected from the prescribed velocity field for which $\text{V}_{1}=\text{V}_{2}$. The G-POD modes contain information on the most dominant features of the flow. From these modes we can see that some interesting dynamics occur in the region where the flow changes direction. The second and the third mode also seem to be quite similar, but slightly shifted. This is common for advection dominated flows and results in a slow singular value decay [5]. For the space-local approaches we find that the basis functions contain less information about the flow, but seem more similar to a generic basis. This likely allows them to generalize better on the validation set. We provide more insight on this in B. For LO-POD we find that the basis functions smoothly decay to zero at the edge of the subdomain, as enforced by the kernel introduced in A. This allows LO-POD to produce smooth approximations. The basis resulting from L-POD does not guarantee this.
@@ 5.6.2 ROM performance
In this section we discuss the performance of different ROMs and compare to both the FOM and a finite volume discretization which is twice as coarse as the FOM in each direction ($128\times 128$ as compared to $256\times 256$ for the FOM). This will be referred to as coarse FOM. The results are summarized in Table 1 for a full simulation up to $t=40$ (including the training, validation, and extrapolation region).
Table 1: Computation times (comp time) in seconds on a standard laptop CPU and average solution errors (sol err) computed for the entire simulation $t\in[0,40]$ for each of the ROMs for the 2D advection test case. A coarse FOM with a resolution of $128\times 128$ is also included for comparison. For the space-local approaches the number of DOF are reported as the multiplication $I\times q=r$. Reported computation times are an average of $5$ replica simulations. The average error for the gradient of the solution (grad err) is also depicted.
DOF ($I\times q$) comp time $||\text{sol err}||_{2}$ $||\text{grad err}||_{2}$ $\Delta t$
@@
L-POD $12808^{2}\times 20=$ 1280 0.20700 0.0353 0.213 0.200
LO-POD $9608^{2}\times 15=$ 960 1.02000 0.0354 0.164 0.050
coarse FOM $16384128^{2}=$ 16384 4.57000 0.2260 0.514 0.100
FOM $65536256^{2}=$ 65536 25.9000 - - 0.025
@@
$\text{gradient error}:=\frac{\mathbf{G}\mathbf{u}^{\text{ROM}}_{r}(t)-\mathbf{G}\mathbf{u}^{\text{FOM}}(t)}{||\mathbf{G}\mathbf{u}^{\text{FOM}}(t)||_{2}},$ (44)
where $\mathbf{G}\in\mathbb{R}^{2N}$ is a forward difference approximation of the gradient operator, such that the elements of $\mathbf{G}\mathbf{u}_{r}^{\text{ROM}}$ approximate $\nabla u$ in different parts of the domain. These large errors are likely the result of the limited generalization capabilities of the G-POD basis. For the space-local approaches, we observe that the computation time is roughly two order of magnitude larger than for G-POD. This is due to the fact that the basis is larger. However, using the space-local approaches we still obtain a speedup between one and two orders of magnitude with respect to the FOM. The L-POD ROM is roughly 5 times faster than the LO-POD ROM. This is mostly due to the fact that the explicit RK4 scheme (fourth order accurate) used for L-POD allows for larger time steps without loss of accuracy, as compared to the Crank-Nicolson scheme (only second order accurate) used for LO-POD, see C for an analysis on the time step size. L-POD and LO-POD both produce a similar solution error. However, looking at the gradient error LO-POD obtains a lower error. This is likely due to the fact that LO-POD does not suffer from discontinuities at the subdomain boundaries, while L-POD does, see Figure 2. Both space-local approaches outperform the coarse grid finite volume discretization both in computation time and error metrics.
@@
We find that the G-POD ROM performs well within the training region, but it does not extrapolate well. For the space-local approaches this is not an issue as both perform comparably, both inside and outside the training region, and for both the solution error and the energy. These approaches also outperform the coarse FOM, which slowly accumulates error during the simulation.

### v3

@@ 5.2 Projection error
For the construction of the L-POD and LO-POD ROMs, we have to determine the number of subdomains $I$ and the number of modes per subdomain $q$, with $Iq=r$, which results in a basis that generalizes best outside the training data. To do this, we evaluate the projection error, see (41), on both the training and validation data. The results are depicted in Figure 6 for different $I$, averaged over both the training (solid line) and validation data (dashed line).
Figure 6: (Top) Projection error evaluated over the training (solid line) and validation (dashed line) data for different II for the space-local approaches. (Bottom) Projection error for $I=10$ for L-POD and LO-POD evaluated over the training (solid line) and validation (dashed line) data. The projection error for G-POD is also depicted.
A basis that generalizes well should perform well on both the training and validation data. We observe that for $I=5$, the training error is lowest, but the validation error is rather large, indicating that the basis does not generalize well. For higher values of $I$, the training and validation errors are much closer, but tend to increase with increasing $I$. In this case, $I=10$ is considered optimal, as the training and validation errors are small and of similar size. This is true for both L-POD and LO-POD. For the remainder of this text, we will therefore stick to $I=10$ for the construction of the local ROMs. In general, the choice of $I$ is likely problem-dependent, and an a priori study similar to Figure 6 needs to be performed to determine an optimal value.
Next, we compare the performance of the space-local approaches against each other, as well as to G-POD. This is also displayed in Figure 6 (bottom plot). We observe that the projection error converges faster for LO-POD, as compared to L-POD. This can be explained by the fact that the LO-POD basis functions are constructed from twice as many finite difference points as for L-POD, due to the overlapping subdomains. This means the fit is carried out over twice as many DOF, which yields a lower projection error. The difference between convergence on the training and validation data is slight for both space-local approaches. On the other hand, for G-POD, the convergence of the projection error on the training data is very fast, see (13). However, on the validation data, the error hardly converges. This is because the G-POD is only suited for representing the wave on the left side of the domain, as detailed in the next section. We conclude that the space-local approaches yield a basis that generalizes better than the global approach for this particular problem.
@@ 5.3 Resulting basis
Figure 7: The first six G-POD modes and the first six space-local POD modes for L-POD and LO-POD. For visualization purposes, the space-local POD modes are displayed on adjacent subdomains.
We find that the G-POD modes are only nonzero where the traveling wave is represented in the training data. This explains why the error on the validation data hardly converged in Figure 6. For the space-local approaches, where a common basis is “copied/pasted” across the entire domain, this is not an issue. Regarding L-POD, we observe that the obtained basis functions do not smoothly decay to zero at the edge of the subdomain, but instead end abruptly. Finally, we observe that for LO-POD the post-processing procedure in (28) indeed results in a basis that smoothly decays on the edge of the subdomains.
@@ 5.4 Accuracy and energy conservation of ROM
Having obtained a basis for each of the POD approaches we construct a set of ROMs. Based on the results in Section 5.2, we select $r=60$, $I=10$, and $q=6$ for the space-local approaches. To keep the size of the basis the same, we choose $r=60$ for G-POD. The ROM results are obtained by evaluating (35) from the initial condition in (40). The resulting simulations up to $t=5$ are presented in Figure 8.
@@
For G-POD we find that after the training data the performance degrades significantly. This exposes the limitations of G-POD, as it is unable to adapt to wave traveling outside the training range. The space-local approaches perform better in this regard. Both L-POD and LO-POD are capable of extrapolating the traveling of the wave past the training region. However, L-POD seems to suffer from discontinuities in $\mathbf{u}_{r}^{\text{ROM}}$, as the edges of the subdomains become increasingly visible as the simulation progresses. LO-POD does not suffer from this issue and smoothly extrapolates the solution past the training region. This means the smooth reconstruction coming from LO-POD indeed improves the quality of the simulation for the same number of DOF.
@@
The results are depicted for both the FOM time step size $\Delta t$ (solid line) and a five times larger time step $5\cdot\Delta t$ (dashed line). This larger time step is based on an eigenvalue analysis of the ROM operator, also presented in Figure 9. For this analysis, we determined the eigenvalues of the operator $\mathbf{D}$ projected on the ROM basis. This operator is given by $\mathbf{P}\mathbf{D}\mathbf{P}$. It can be shown that replacing $\mathbf{D}$ in the FOM, see (6), by this projected operator is equivalent to integrating the ROM, see (35). The eigenvalues $\lambda_{i}$ are ordered according to the magnitude of their absolute values. We observe that the largest eigenvalue for the space-local approaches is roughly five times smaller than that of G-POD and the FOM. This is likely due to the fact that the G-POD basis functions contain higher frequencies than the space-local basis functions, see Figure 7. Using a smaller G-POD basis could alleviate this. Based on this analysis, we also evaluate the performance of the ROMs for a five times larger time step [eigenvalues_BALDAUF20086638].
@@
Looking at the solution error, we observe that G-POD performs best in the training region (the error is so small that it is outside the plotting range). However, after leaving the training region ($t>1$), the performance degrades rapidly. On the other hand, for the space-local approaches, we do not observe this jump of error outside the training region, but rather a steady increase. Importantly, outside the training region, the space-local approaches are much more accurate than G-POD. In particular, LO-POD improves by more than one order of magnitude upon both G-POD and L-POD outside the training region.
@@
When increasing the time step size by fivefold, the simulation quickly becomes unstable for G-POD. For L-POD, both time steps yield stable simulations, giving results that are very close to each other. For LO-POD and a smaller time step size, the error seems to converge to an equilibrium. However, for a larger time step, it increases steadily. This means there is likely still a benefit to taking a smaller time step. An interesting continuation of this research would be to find a way to systematically determine the “sweet spot” between increasing the time step and maintaining accuracy of the ROM. This could for example, be done by considering several different time step sizes and both implicit and explicit integration methods, or adaptive time-stepping based on an error estimator [Johnson1987].
@@
In terms of energy conservation, we observe that all ROMs conserve the energy as predicted by the theoretical analysis, except for a time discretization error incurred by the use of RK4. This error increases when the time step size is increased. Only for G-POD with an increased time step size, the simulation becomes unstable. For LO-POD, the numerical error in the instantaneous change in energy, see (39), is the largest. This is likely due to the linear system that needs to be solved to evaluate the ROM for a non-orthogonal basis, see (35). However, the time discretization error is the primary source of error, as the change in energy during the simulation is roughly the same for L-POD and LO-POD.
@@ 5.5 Convergence with increasing ROM dimension
Next, we look at the convergence of the solution error as we increase the size of ROM basis $r$. In addition, we compare the ROMs to a finite difference discretization with the same number of DOF. Note that for the ROMs $\text{DOF}=r$. For the finite difference simulations, we first project the solution on the FOM grid using linear interpolation. We then compute the difference, such as in (41), and compute the $L_{2}$-norm to quantify the error. For the local ROMs we stick to $I=10$ subdomains, while increasing the number of basis functions per subdomain $q$ to increase $r$. We consider a maximum of $r=100$ for the ROMs.
@@
Regarding the local ROMs, we observe rapid convergence of the solution error as we increase the number of DOF, with LO-POD consistently outperforming L-POD. This is in line with the projection error convergence discussed in Section 5.2. Regarding G-POD, we observe no convergence of the solution error. Regarding the finite difference discretization, we find it converges at a much slower rate than the space-local ROMs.
To obtain a conclusive answer about the scaling of the ROMs, we considered the number of nonzero entries in the ROM operator $\mathbf{A}$. This is also depicted in Figure 10. Here we find that the number of nonzero entries scales according to a power law, with L-POD scaling the most favorably. This is in line with the discussion presented in Section 4.2. Hyper-reduction can be carried out to further reduce the computational cost of the ROMs [hyperreduction_1, hyperreduction_2, KLEIN2024112697]. Finally, one could also increase the time step size, as discussed in Section 5.4, to further decrease the computational cost of the space-local ROMs.
5.6 2D advection-diffusion equation
To further evaluate the viability of the space-local ROMs, we consider the 2D advection-diffusion equation:
@@
where some quantity $u(\mathbf{x},t)\in\mathbb{R}$ is advected through space $\mathbf{x}=\begin{bmatrix}x&y\end{bmatrix}^{T}\in\mathbb{R}^{2}$ by a stationary velocity field $\mathbf{V}\in\mathbb{R}^{2}$ which is divergence free, i.e. $\nabla\cdot\mathbf{V}=0$. The fact that the velocity field is now space-dependent, as opposed to the 1D case where it was uniform, adds to the difficulty of the test case. As opposed to the 1D case this system also contains diffusion. This results in a passive spread of $u$ throughout the domain. The diffusion rate is determined by the scalar viscosity $\nu\geq 0$. In our case, we employ a finite volume discretization with a staggered grid for the velocity field [benjamin_thesis]. This is done to preserve the skew-symmetry of the operator. The diffusive term is discretized using a simple second-order scheme.
For the FOM we consider the following setting: A periodic domain $\Omega=[-\pi,\pi]\times[-\pi,\pi]$, a velocity field given by $\mathbf{V}=\begin{bmatrix}\cos(x-y)&\cos(x-y)\end{bmatrix}^{T}$ discretized on a $256\times 256$ uniform grid, and a viscosity of $\nu=10^{-3}$. The initial condition is given by the sum of three Gaussians, each centered on one of the streamlines of the flow where the velocity is highest:
$\begin{split}u(\mathbf{x},0)=&-\exp(-2x^{2}-2y^{2})+\exp(-2\left(x-\frac{\pi}{2}\right)^{2}-2\left(y+\frac{\pi}{2}\right)^{2})\\ &+\exp(-2\left(x+\frac{\pi}{2}\right)^{2}-2\left(y-\frac{\pi}{2}\right)^{2}).\end{split}$ (43)
The FOM is integrated in time using a RK4 scheme with $\Delta t=0.025$. Every 16 time steps, we save a snapshot of $u(x,t)$ for the construction of the ROM basis. Note that the performance of the ROMs converge to the FOM, when doing the Galerkin projection. To resolve this closure, models can be included in the ROM [snyder2022reducedordermodelclosures]. The snapshots collected in the interval $t=[0,12]$ are used as training data, the interval $t=(16,20]$ is used as validation data, and the interval $t=(20,40]$ is used to test the extrapolation capabilities of the ROMs. The ROMs are also integrated in time explicitly using a RK4 scheme. However, for the LO-POD ROM we resort to the implicit Crank-Nicolson method, see [FEM_book], to limit the computational cost. This scheme enables us to achieve second-order accuracy in time, while solving only a single linear system at each time step, rather than at each evaluation of (35) in the RK4 scheme. To solve the resulting system efficiently, we employ an LU-decomposition [MITTAL2002131_LU]. This is required, as the LO-POD basis is non-orthogonal, i.e. $\mathbf{S}\neq\mathbf{I}$.
Regarding the subdomain decomposition, we use the fact that the FOM is discretized on a uniform grid. In this way, we can simply subdivide the domain uniformly into subdomains. For L-POD, this results in a straightforward decomposition. For LO-POD the decomposition is chosen such that each subdomain overlaps with its four neighboring subdomains which share a vertex in the middle of the considered domain. This is depicted in Figure 11.
@@
The number of subdomains $I$ for the space-local approaches and the number of modes $q$ is chosen according to the performance on the validation set. This is presented in B. Based on this, we settle on $I=8\times 8$ and $q=20$ for L-POD and $q=15$ for LO-POD. We settle on these values for $q$ as these are the lowest values which surpass a projection error threshold of $10^{-2}$ on the validation set for this value of $I$. The value of $I$ is chosen as it generalizes well from the training data set to the validation data set. For G-POD, we choose $r=q=31$, as the snapshot matrix contains $31$ snapshots for G-POD such that the SVD only produces $31$ basis functions. For the size of the time step, we choose the largest value that does not degrade the ROM performance. This is presented in C. The time step sizes used are presented in Table 1.
@@
The resulting basis computed with the SVD for each of the POD approaches is depicted in Figure 12.
@@
From the basis functions, we can see that the flow runs diagonally across the domain. This is expected from the prescribed velocity field for which $\text{V}_{1}=\text{V}_{2}$. The G-POD modes contain information on the most dominant features of the flow. From these modes, we can see that some interesting dynamics occur in the region where the flow changes direction. The second and the third mode also seem to be quite similar, but slightly shifted. This is common for advection-dominated flows and results in a slow singular value decay [brunton2019data]. For the space-local approaches, we find that the basis functions contain less information about the flow, but seem more similar to a generic basis. This likely allows them to generalize better on the validation set. We provide more insight on this in B. For LO-POD we find that the basis functions smoothly decay to zero at the edge of the subdomain, as enforced by the kernel introduced in A. This allows LO-POD to produce smooth approximations. The basis resulting from L-POD does not guarantee this.
@@ 5.6.2 ROM performance
In this section, we discuss the performance of different ROMs and compare to both the FOM and a finite volume discretization, which is twice as coarse as the FOM in each direction ($128\times 128$ as compared to $256\times 256$ for the FOM). This will be referred to as coarse FOM. The results are summarized in Table 1 for a full simulation up to $t=40$ (including the training, validation, and extrapolation region).
Table 1: Computation times (comp time) in seconds on a standard laptop CPU and average solution errors (sol err) computed for the entire simulation $t\in[0,40]$ for each of the ROMs for the 2D advection test case. A coarse FOM with a resolution of $128\times 128$ is also included for comparison. For the space-local approaches, the number of DOF is reported as the multiplication $I\times q=r$. Reported computation times are an average of 55 replica simulations. The average error for the gradient of the solution (grad err) is also depicted.
DOF ($I\times q$) comp time $\text{sol err}||_{2}$ $\text{grad err}||_{2}$ $\Delta t$
@@
L-POD $8^{2}\times 20=1280$ 0.20700 0.0353 0.213 0.200
LO-POD $8^{2}\times 15=960$ 1.02000 0.0354 0.164 0.050
coarse FOM $128^{2}=16384$ 4.57000 0.2260 0.514 0.100
FOM $256^{2}=65536$ 25.9000 - - 0.025
@@
[…]
where $\mathbf{G}\in\mathbb{R}^{2N\times N}$ is a forward difference approximation of the gradient operator, such that the elements of $\mathbf{G}\mathbf{u}_{r}^{\text{ROM}}$ approximate $\nabla u$ in different parts of the domain. These large errors are likely the result of the G-POD basis’s limited generalization capabilities. For the space-local approaches, we observe that the computation time is roughly two orders of magnitude larger than for G-POD. As we have access to more space-local POD modes than global ones, we choose to include enough to accurately represent the solution, see B for details. This increases the computation time with respect to G-POD, see Table 1 for the exact number of DOF for each ROM. However, using the space-local approaches, we still obtain a speedup between one and two orders of magnitude with respect to the FOM. The L-POD ROM is roughly 5 times faster than the LO-POD ROM. This is mostly because the explicit RK4 scheme (fourth order accurate) used for L-POD allows for larger time steps without loss of accuracy, as compared to the Crank-Nicolson scheme (only second order accurate) used for LO-POD, see C for an analysis on the time step size. As mentioned earlier, the Crank-Nicolson scheme is used for LO-POD as it limits the number of linear systems needed to be solved per time step to one. L-POD and LO-POD both produce a similar solution error. However, looking at the gradient error LO-POD obtains a lower error. This is likely because LO-POD does not suffer from discontinuities at the subdomain boundaries, while L-POD does, see Figure 3. Both space-local approaches outperform the coarse grid finite volume discretization, both in computation time and error metrics.
@@
We find that the G-POD ROM performs well within the training region, but it does not extrapolate well. For the space-local approaches, this is not an issue as both perform comparably, both inside and outside the training region, and for both the solution error and the energy. These approaches also outperform the coarse FOM, which slowly accumulates error during the simulation.
5.7 Applicability to 2D Navier-Stokes
As a final test case, we evaluate the applicability to the 2D incompressible Navier-Stokes equations by assessing how well the POD bases generalize to representing the solution. The incompressible Navier-Stokes equations read:
$\frac{\partial\mathbf{V}}{\partial t}+(\mathbf{V}\cdot\nabla)\mathbf{V}$ $=-\frac{1}{\rho}\nabla p+\nu\Delta\mathbf{V}+\mathbf{f},$ (45a)
$\nabla\cdot\mathbf{V}$ $=0,$ (45b)
for a 2D velocity field $\mathbf{V}(\mathbf{x},t)\in\mathbb{R}^{2}$. For the density we choose $\rho=1$ and for the kinematic viscosity $\nu=\frac{1}{1000}$. For the forcing we take $\mathbf{f}(\mathbf{V},\mathbf{x},t)=\begin{bmatrix}\sin(4y)&0\end{bmatrix}^{T}-0.1\mathbf{V}$, as to simulate Kolmogorov flow. This setup is often used for evaluating data-driven turbulence modeling strategies [Kochkov_2021, shankar2024differentiableturbulenceclosurepartial, Agdestein2025MLLES]. The PDE is solved numerically using the second-order accurate scheme presented in [harlow1965numerical] on a periodic domain $\Omega=[-\pi,\pi]\times[-\pi,\pi]$. The computational grid consists of $2048\times 2048$ grid cells. The simulation is initialized with energy content in the large wave-numbers, and a warm-up period of 25 time units is employed before collecting the training data. After the warm-up period, training data is collected on the interval $t=[0,10]$ and validation data on the interval $t=[100,110]$, to remove any temporal correlation. Both the validation and training data consist of 50 snapshots of the vorticity $\boldsymbol{\omega}=\nabla\times\mathbf{V}\in\mathbb{R}^{3}$. For 2D, only the third component of this vector is nonzero, which we shall refer to simply as $\omega$. We choose to represent the vorticity as for periodic domains in 2D, all the information of the flow is contained within this scalar field [lequeurre2018vorticitystreamfunctionformulations].
To begin our analysis, we examine the projection error for both the training and validation datasets using an $8\times 8$ decomposition of subdomains, which is identical to the 2D advection-diffusion case, see Figure 15.
Figure 15: Projection error evaluated over the training (left) and validation (right) data set for the 2D Navier-Stokes test case. $I=1\times 1$ corresponds to G-POD. Solid dots correspond to L-POD, stars to LO-POD.
Similar to the previous test cases, we find that the projection error of the G-POD converges fast on the training set. However, upon examining the validation set, we find that the space-local approaches generalize significantly better.
In Figure 16, we display the reconstruction of the first snapshot of the validation set produced by the different bases.
Figure 16: First vorticity snapshot of the validation data projected on the different POD bases with varying number of modes per subdomain $q$ for the space-local approaches.
We depict the reconstructions for both $q=10$ and $q=50$ modes per subdomain for the space-local approaches, and all the available POD modes, i.e. $r=50$, for G-POD. We find that the space-local approaches capture the flow characteristics quite well, with $8\times 8\times 50=3200$ DOF, whereas G-POD does not capture any of the characteristics. Regarding L-POD, we find that using a lower number of modes per subdomain $q$ worsens the discontinuities at the subdomain boundaries, whereas LO-POD suffers from noisier reconstructions for low $q$.